You are contemplating your retirement needs. You want to have an annual retirement income of $70,000 for 35 years of retirement and expect an interest rates of 2% over your retirement. You expect to retire 40 years from now. How much would you need to put into your RRSP annually starting one year from now for the next 40 years to achieve your retirement financial goal assuming you can get 3% on your investments until retirement? Round your answer to the nearest dollar.
Answer #1

Value of retirement income after 40 years from now (at t = 40) = $70,000*PVIFA(2%,35)

= $70,000*24.9986

= $17,49,902

Present value of the above today (at t = 0) = Value of retirement income after 40 years from now (at t = 40)*PVIF(3%,40)

= $17,49,902*0.3066

=$5,36,519.95

=$5,36,520

.

Let the amount to be invested into the RRSP annually starting one year from now for the next 40 years to achieve the retirement financial goal assuming we can get 3% on our investments until retirement be “$x”

Present Value of the above = x*PVIFA(3%,40)

=x*23.1148

=$23.1148x

.

Therefore,

=23.1148x = 5,36,520

Therefore, x = 5,36,520/23.114 = $23,211.10

.

Therefore , the answer is $23,211.10